# Stimulated absorption of single gravitons: First light on quantum gravity

### v1

@@ Abstract
    Is the energy content in the graviton the same as in a photon? In other words, is Planck’s constant that governs the exchange of energy with gravity the same as for all other interactions, $\hbar_{G}=\hbar$? One can for example imagine a natural number $k$ such that $\hbar_{G}=k\hbar$, and thus the energy exchange with matter proceeds as $E=khf$. One would then expect not the transition to the first excited state, but to the k-th excited state. We note that this would not violate any current experimental data. Similarly, for a general $\hbar_{G}\neq\hbar$, one can monitor for seemingly off-resonant excitations for which the frequencies between wave and matter would match to $f_{G}=\frac{\hbar}{\hbar_{G}}f$.
@@
    Is the probability of stimulated absorption the same as the probability of stimulated emission? This would amount to testing the gravitational Einstein coefficients, and whether $B_{12}^{(G)}=B_{21}^{(G)}$. Such a test would require the preparation of the resonator not just in the ground state, but also in the first excited state, monitoring quantum jumps to the ground state due to stimulated emission of gravitons.
@@
    Does the absorption of gravitons follow the quadrupole-formula as expected? Making use of independent LIGO detections, one can test if the probability of single graviton events is indeed as expected from a quadrupolar interaction with the collective mode, as given by equation (3). This would provide experimental input on possible scalar or vector components of the interaction at the quantum level.
@@
    Do gravitational waves also carry quantized momentum $p=hf/c$? This, for example, is central to quantum noise from gravitons in interferometers [41], but such fluctuations remain very difficult to observe. However, if such observations of graviton noise or imprints of gravitons on the CMB [42] could be made, in combination with tests of the $E=hf$ relation as we proposed, they would firmly establish a particle picture as they test complementary aspects of quantization. Together such tests could therefore provide a bona-fide test of quantum gravity.
@@
This list is by no means exhaustive. For example, one can also consider tests of the wave-particle duality through statistical properties, as evident in Einstein’s early works [27, 43, 30]. These examples show the potential of a single-graviton-detector as we proposed for fundamental explorations of the interaction between matter and gravity at the single graviton level. Thus, while such experiments with stimulated emission cannot serve as definite proofs, they clearly provide a wide range of opportunities to test aspects of quantum gravity. While we do not anticipate deviations from expected physics, it would provide empirical experimental evidence where none exists so far, and there is always room for surprises (sometimes leading to unexpected explanations for other open problems in physics). Furthermore, together with tests of gravitational shot noise from fluctuations in interferometers or the CMB, they would rule out most semi-classical models which otherwise might explain each effect in isolation, and thus can provide very strong evidence of quantization.
@@
Conclusions. We conclude this essay with a positive outlook on experimental tests of quantum gravity. Similar to the development of the early quantum theory of the 20th century, the mid 21st century will likely see first experimental demonstrations of the quantized interaction between gravity and matter, such as through our proposed single graviton exchange [3]. Other experimental proposals exist as well that probe other aspects of quantum features, such as testing how quantum superpositions source Newtonian gravity and cause entanglement [44, 45, 46, 47], or testing speculative phenomenological models [48, 49, 50, 51]. Here we have argued that a single graviton detector would provide a complementary approach and offer a range of opportunities to test the very basic principles of quantized interaction between gravitational waves and matter, in close analogy to historic experiments that established the existence of the photon in the early 20th century. We are thus hopeful that there are now realistic paths for experimental input to quantum gravity from laboratory experiments.

### v2

@@ IV Tests with single-graviton interactions
    Is the energy content in the graviton the same as in a photon? In other words, is Planck’s constant that governs the exchange of energy with gravity the same as for all other interactions, $\hbar_{G}=\hbar$? One can for example imagine a natural number $k$ such that $\hbar_{G}=k\hbar$, and thus the energy exchange with matter proceeds as $E=khf$. One would then expect not the transition to the first excited state, but to the k-th excited state. Similarly, for a general $\hbar_{G}\neq\hbar$, one can monitor for seemingly off-resonant excitations for which the frequencies between wave and matter would match to $f_{G}=\frac{\hbar}{\hbar_{G}}f$. We note that this would not violate any current experimental data on classical gravitational wave detection, and neither from other non-gravitational parts of quantum physics, as the standard constant $\hbar$ is unaltered in the non-gravitational sector. The reasoning for probing whether $\hbar_{G}=\hbar$ is to test whether the individual quanta of gravity are of the same nature as the quanta of light and matter.
@@
    Is the probability of stimulated absorption the same as the probability of stimulated emission? This would amount to testing the gravitational Einstein coefficients, and whether $B_{12}^{(G)}=B_{21}^{(G)}$. Such a test would require the preparation of the resonator not just in the ground state, but also in the first excited state, monitoring quantum jumps to the ground state due to stimulated emission of gravitons. Since both stimulated emission and absorption are both accessible with the type of detectors we propose, this test would provide a window into the nature of energy conservation for the exchange of individual quanta.
@@
    Does the absorption of gravitons follow the quadrupole-formula as expected? Making use of independent LIGO detections, one can test if the probability of single graviton events is indeed as expected from a quadrupolar interaction with the collective mode, as given by equation (3). This would provide experimental input on possible scalar or vector components of the interaction at the quantum level, or otherwise confirm the spin-2 nature of the particle.
@@
    Do gravitational waves also carry quantized momentum $p=hf/c$? This, for example, is central to quantum noise from gravitons in interferometers [52], but such fluctuations remain very difficult to observe. However, if such observations of graviton noise or imprints of gravitons on the CMB [53] could be made, in combination with tests of the $E=hf$ relation as we proposed, they would firmly establish a particle picture as they test complementary aspects of quantization, analogously to photons [17, 34]. Together such tests could therefore provide a bona-fide test of quantum gravity.
Figure 2: Detection of single gravitons from passing waves by stimulated processes can provide experimental evidence for linearized quantum gravity. The five proposed tests in the main text are illustrated here: (i) Probing the energy content in a single quantum of gravitational energy, (ii) testing universality of the interaction, (iii) comparing stimulated emission and absorption, (iv) testing the spin of a single graviton, (v) probing the quantization of momentum.
@@
This list is by no means exhaustive. For example, one can also consider tests of the wave-particle duality through statistical properties, as evident in Einstein’s early works [30, 54, 33]. These examples show the potential of a single-graviton-detector as we proposed for fundamental explorations of the interaction between matter and gravity at the single graviton level. They can highlight some of the very fundamental properties that any broader theory of quantum gravity would have to obey, offering empirical anchoring of some of the basic principles. Importantly, these tests require only access to stimulated processes, and are thus attainable once single graviton absorption can be achieved. The additional requirements are the ability to detect various energy levels and modes in the resonator, prepare different detector materials and energy eigenstates as initial states, and to be able to correlate to classical detections via LIGO or other means. Thus, while such experiments with stimulated processes cannot serve as definite proofs, they clearly provide a wide range of opportunities to test aspects of quantum gravity. While we do not anticipate deviations from expected physics, these tests would provide empirical experimental evidence where none exists so far, and there is always room for surprises (sometimes leading to unexpected explanations for other open problems in physics). Furthermore, together with tests of gravitational shot noise from fluctuations in interferometers or the CMB, they could rule out most semi-classical models which otherwise might explain each effect in isolation, and thus can provide very strong evidence of quantization.
V Conclusions
@@
We conclude with a positive outlook on experimental tests of quantum gravity. Similar to the development of the early quantum theory of the 20th century, the mid 21st century will likely see first experimental demonstrations of the quantized interaction between gravity and matter, such as through our proposed single graviton exchange [3]. Other experimental proposals exist as well that probe other aspects of quantum features, such as testing how quantum superpositions source Newtonian gravity and cause entanglement [55, 56, 57, 58], or testing speculative phenomenological models [59, 60, 61, 62]. Here we have argued that a single graviton detector would provide a complementary approach and offer a range of opportunities to test the very basic principles of quantized interaction between gravitational waves and matter, in close analogy to historic experiments that established the existence of the photon in the early 20th century. Combined with other tests, the quantum nature of gravity can be probed from a variety of perspectives. We are thus hopeful that there are now realistic paths for experimental input to quantum gravity from laboratory experiments.
@@ Acknowledgements.
Appendix A Detector response at finite duration
In classical gravitational wave literature [63, 64, 65], a resonant bar detector will be modeled with damping
$\ddot{\xi}(t)+\gamma\dot{\xi}(t)+\omega^{2}\xi(t)=F_{G}(t)/M_{eff}$ (5)
where $\xi(t)$ represents the mass’ perturbation from equilibrium position, $\omega$ the resonance of the detector, $\gamma=\omega/Q$ the damping as a factor of the quality factor $Q$, $M_{eff}$ the effective mass of the mode, and $F_{G}(t)$ the effective gravitational tidal force due to the incident gravitational wave. For $F_{G}(t)=-2Lh\nu^{2}/\pi^{2}\ \textrm{Re}\left[e^{-i\nu t}\right]$, this produces the solution [65]:
$\xi(t)=\,\frac{2Lh\nu^{2}}{\pi^{2}}\frac{\left(\nu^{2}-\omega^{2}\right)\cos(\nu t)-\gamma\nu\sin(\nu t)}{\left(\nu^{2}-\omega^{2}\right)^{2}+\gamma^{2}\nu^{2}}\,.$ (6)
This result might seem problematic for the resonant $\nu=\omega$, non-damping $\gamma=0$ case, however there is a specific solution for the exact resonance without damping [65]
$\xi_{res}(t)=\frac{2Lh\omega t}{\pi^{2}}\sin(\omega t)\,.$ (7)
A perfect detector will keep absorbing energy from the gravitational wave ad infinitum, or in practice, until the damping of the detector becomes a significant factor as the system reaches a semi-stable state [65, 63].
However, for the binary mergers detected by LIGO, the duration of the GW in any particular frequency range is very short, and chirps upwards in frequency until the merger is finalized [1, 2]. The incident gravitational wave persists in the resonant band of our detector only for some short time $\tau$, and thus the solution (7) is valid for time $\tau$ without damping playing the limiting role. Applying this time scale to the resonant result, we can calculate the expected graviton excitation probability using the relation $\frac{dP}{dt}=\frac{1}{\hbar\omega}\frac{dE}{dt}$ [64]:
$P_{0\rightarrow 1}\simeq\frac{E}{\hbar\omega}=\frac{M_{eff}}{2\hbar\omega}\Bigl(\left\langle\dot{\xi}(t)^{2}\right\rangle+\omega^{2}\left\langle\xi(t)^{2}\right\rangle\Bigr)=\frac{h^{2}ML^{2}\omega^{3}\tau^{2}}{\hbar\pi^{4}}$ (8)
where $M_{eff}=M/2$ for a simple bar. This result agrees with the quantum excitation probability computed in Ref. [3] and given in eq. (3) within the finite duration window $\tau$.